OpenStudy (anonymous):

Find the slope of the line whose equation is 6x - 3y = 12.

OpenStudy (help!!!!):

isolate the y on one side

OpenStudy (help!!!!):

so put it into the y-intercept form

OpenStudy (help!!!!):

y=mx+b

OpenStudy (anonymous):

so slope is 2

OpenStudy (anonymous):

-3y=12-6x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now isolate the y

OpenStudy (help!!!!):

now divide by -3

OpenStudy (anonymous):

y=-4-2x